OpenStudy (anonymous):

What is the slope of the line given by the equation 7y + 5x = 2 (BEST ANSWER GETS MEDAL!)

OpenStudy (texaschic101):

the first thing you need to do is put your equation in y = mx + b form. Once that is done, you slope can be found in the m position.

OpenStudy (anonymous):

7 y = - 5x + 2 y = -5/7 x = 2/7 therefore the slope is -5/7

OpenStudy (anonymous):

Thank you! :)